OpenStudy (anonymous):

How do you right a equation of a line that passes through the points of (0,-7) and a slope of -2

OpenStudy (anonymous):

y+7=-2(x-0)

OpenStudy (anonymous):

y+7=-2x+2

OpenStudy (anonymous):

y=-2x-5